OpenStudy (anonymous):

using wallis formula solve: integrate from negative pi over two to pi over two. sin^3 t cos^10 t dt

OpenStudy (shubhamsrg):

what is wallis formula ?

OpenStudy (shubhamsrg):

nevermind, just substitute cost = u see if it helps.

OpenStudy (anonymous):

I have seen Wallis Formula before, unfortunately never in combination with a product of sin and cosine. Hence I would solve this integral using a substitution just as suggested: \[\Large \int \sin^2(x)\sin(x)\cos^{10}(x)dx \\ \Large \int(1-\cos^2(x))\sin(x)\cos^{10}(x)dx \\ \\ \Large \int \sin(x)\cos^{10}(x)dx-\int\sin(x)\cos^{12}(x)dx \] But since this answer isn't sufficient enough for this kind of problem, I will see if I can find again my notes/sites on Wallis Formula.